A license plate number consists of three letters followed by three digits. How many different license numbers can be formed if repetition of letters and digits is allowed? How many different license numbers can be formed if repetition of letters and digits is not allowed?
Answer by checkley77(12844)

26*26*26*10*10*10=26^3*10^3=17,576*1,000=17,576,000 ans. with repeats.
26*25*24*10*9*8=15,600*720=11,232,000 ans. without repeats.